After completing a long and successful career as senior vice president for a large bank, you are preparing for retirement. After visiting the human resources office, you have found that you have several retirement options to choose from:

  1. An immediate cash payment of $1.18 million.
  2. Payment of $68,000 per year for life.
  3. Payment of $58,000 per year for 4 years and then $78,000 per year for life (this option is intended to give you some protection against inflation).


You believe you can earn 9 percent on your investments and your remaining life expectancy is 8 years.

Required:
1.
Calculate the present value of each option. (Future Value of $1, Present Value of $1, Future Value Annuity of $1, Present Value Annuity of $1.) (Use appropriate factor(s) from the tables provided. Do not round intermediate calculations. Enter your answers in whole dollars, not in millions. Round the final answer to nearest whole dollar.)
Answer #1

PV of cash Inflow

1. Immediate cash payment= 1.18 million

2. Payment of $58000 per year for life

PV of annuity= 5.53

PV of cash inflow= 5.53 x 58000= 321020

3. Payment of 58000 for first 4 years, 78000 for 5th - 8th year

PV of cash inflows for year 1 -4= 3.239 x 58000= 187904

Pv of 5th year cash flow= .649 x 78000= 50695

PV of 6th year cash flow = .596 x 78000= 46509

PV of 7th year cash flow= .547 x 78000= 42669

PV of 8th year cash flow= .501 x 78000 = 39146

PV of total cash inflow= 187904+50695+46509+42669+39146= 366923